a tennis ball “A” is released from rest down a 10.0 m long inclined ramp with a uniform acceleration of 5.0 m/s2. Another tennis ball “B” is initially located at the same height as ball “A” right above the lower edge of the ramp. Ball “B” is thrown upward with some initial speed at the same instant as the release of ball “A”.


1
Expert's answer
2021-07-12T02:59:02-0400

Gives

Ramp length (l)=10m

a=5m/sec2



"s=lsin\theta"

Newton's third law

"v^2=u^2+2as"

u=0

"s=lsin30=10sin30\u00b0=5m"

"v^2=0+2\times5\times5=50m/sec"

"v=5\sqrt{2}m/sec"